# On the expressiveness and spectral bias of KANs

### v1

@@ Abstract
Kolmogorov-Arnold Networks (KAN) [46] were very recently proposed as a potential alternative to the prevalent architectural backbone of many deep learning models, the multi-layer perceptron (MLP). KANs have seen success in various tasks of AI for science, with their empirical efficiency and accuracy demonstrated in function regression, PDE solving, and many more scientific problems.
@@ 1 Introduction
Recently, in [46], a novel architecture called Kolmogorov-Arnold Networks (KANs) was proposed as a potentially more accurate and interpretable alternative to standard multi-layer perceptron (MLP) [18, 29]. The KAN architecture leverages the Kolmogorov-Arnold representation theorem (KART) [35, 36] to parameterize functions. KANs share the same fully connected structures as MLPs, while putting learnable activation functions on edges as opposed to a fixed activation function on nodes for MLPs. B-splines are used to parameterize the learned nonlinearity and in practice one can go beyond the two-layer construction indicated by KART. KANs can be conceptualized as a hybrid of splines and MLPs, and the combination of compositional nonlinearity with 1D splines contributes to both the accuracy and interpretability of KANs.
@@
In this article, we study the approximation theory and spectral bias of KANs and compare them with MLPs. Specifically, we show that any MLP with the ReLUk activation function can be reparameterized as a KAN with a comparable number of parameters. This establishes that the representation and approximation power of KANs is at least as great as that of MLPs. On the other hand, we also show that any KAN (without SiLU non-linearity) can be represented using an MLP. However, the number of parameters in the MLP representation is larger by a factor proportional to the grid size of the KAN. While we do not know if our construction is optimal, this suggests that KANs with a large grid size may be more efficient at parameterizing certain classes of functions than MLPs. We combine these results with existing optimal approximation rates for MLPs to obtain approximation rates for KANs on Sobolev spaces.
@@
Next, we study the spectral bias phenomenon in the training of KANs. Standard MLPs with ReLU activations (or even tanh) are known to suffer from the spectral bias [53, 75, 76], in the sense that they will fit low-frequency components first. This is in contrast to traditional iterative numerical methods like the Jacobi method which learn high frequencies first [75]. Although the spectral bias acts as a regularizer which improves performance for machine learning applications [53, 84, 51, 87, 23], for scientific computing applications it may be necessary to learn high-frequencies as well. To alleviate the spectral bias, high-frequency information has to be encoded using methods like Fourier feature mapping [66, 69, 10, 49], or one needs to use nonlinear activation functions more similar to traditional methods; see for example the hat activation function [28] which resembles a finite element basis. We theoretically consider a single KAN layer and show that it does not suffer from the spectral bias by analyzing gradient descent for minimizing a least squares objective. Although our analysis is necessarily highly simplified, we argue that it provides some evidence and intuition showing that KANs will have a reduced spectral bias compared with MLPs.
@@
Finally, we study the spectral bias of KANs experimentally on a wide variety of problems, including 1D frequency fitting, fitting samples from a higher-dimensional Gaussian kernel, and solving the Poisson equation with a high-frequency solutions. Based upon the results of our experiments, KANs consistently suffer less from the spectral bias. This also helps explain why KANs are more subject to noise and overfitting, as observed in [60], while grid coarsening (the opposite of grid extension) helps increase the spectral bias and reduces overfitting. Notice that the highest frequency that a single layer of KANs can learn is restricted to the number of grid points we use. The compositional structure associated with the depth also plays a role in learning the different frequencies. We study the effect of depths, widths, and grid sizes of KANs for the spectral bias, and draw conclusions about best practices when choosing KAN hyperparameters.
@@ 2 Prior Work
A theory on the approximation ability of KANs in the function class of compositionally smooth functions was proposed as KAT in [46]. A convergence result independent of the dimension was obtained, leveraging the approximation theory of 1D splines. Compared with the universal approximation theorem of MLPs, KANs take advantage of the intrinsically low-dimensional compositional representation of underlying functions. This result shares an analogy to the rate in generalization error bounds of finite training samples, for a similar space studied for regression problems; see [30, 34], and also specifically for MLPs with ReLU activations [57]. On the other hand, for general Sobolev or Besov spaces, sharp approximation rates have been obtained for ReLU-MLPs (and more generally MLPs with most piecewise polynomial activation functions) [80, 8, 63]. These rates exhibit the curse of dimensionality, which is unavoidable due to the fact that Sobolev and Besov spaces with fixed smoothness are very large in high dimensions.
@@
There are also subsequent works exploring other parametrizations of activation functions in the KAN formulation, including special polynomials [3, 58, 67], rational functions [4], radial basis function [41, 68], Fourier series [73], and wavelets [12, 59]. Active follow-up research focuses on applying KANs to various domains, such as partial differential equations [72, 62, 55] and operator learning [1, 62, 48], graphs [14, 19, 32, 85], time series [70, 25, 74, 24], computer vision [17, 6, 39, 16, 59, 11], and various scientific problems [43, 44, 77, 26, 37, 40, 5, 42, 50, 52]. In the KAN 2.0 paper [45], multiplication was introduced as a built-in modularity of KANs and the connection between KANs and scientific problems was further established in terms of identifying important features, modular structures, and symbolic formulas. In this article, we focus on the original B-spline formulation, one of the reasons being its alignment with continual learning and adaptive learning; see also [55].
@@
The spectral bias of MLPs has been studied both experimentally and theoretically in a variety of works, see for instance [53, 84, 75, 76, 28, 51, 15, 9, 23, 86, 56] and the references therein. This phenomenon is proposed as explaining the regularizing effect of stochastic gradient descent [53, 23]. So far, the spectral bias of KANs has not been studied in detail, and it is this gap that we aim to close in this work.
@@ 3 Representation and Approximation
In this section, we study how to represent ReLUk MLPs using KANs with degree $k$ splines and vice-versa. Using this, we establish approximation rates for KANs from the corresponding results for MLPs.
@@ 3.1 Review of the KAN Architecture
We recall the following definitions of the Kolmogorov-Arnold Network (KAN) architecture introduced in [46]. We refer to [46] for a thorough treatment of all details mentioned here.
@@
The Kolmogorov-Arnold representation theorem (KART) states that any multivariate continuous function on a bounded domain can be represented as a finite composition of univariate continuous functions and addition. Specifically for a continuous $f:[0,1]^{n}\to\mathbb{R}$, there exists continuous 1D functions $\phi_{q,p},\Phi_{q}$ such that (see for instance [13])
@@
Inspired by KART corresponding to a two-layer neural network representation, the authors in [46] unified the inner and outer functions via the proposed KAN layers and generalized the idea of composition to arbitrary depths. A KAN can be represented by an integer array $[n_{0},n_{1},\cdots,n_{L}]$. Here $L$ is the number of layers and $n_{0}$ and $n_{L}$ are input and output dimensions respectively. $n_{l}$ denotes the number of neurons in the $l$-th neuron layer. Usually, we choose $n_{1}=\cdots=n_{L-1}=W$ as the width and refer to $L$ as the depth of the KAN. Layer $l$ maps a vector $\mathbf{x}_{l}\in\mathbb{R}^{n_{l}}$ to $\mathbf{x}_{l+1}\in\mathbb{R}^{n_{l+1}}$, via nonlinear activations and additions, in the matrix form as
@@
A very important structure of KANs proposed in [46] is the grid extension technique, inheriting the multi-level fine-graining from splines. One can first train KANs with fewer parameters to a desired accuracy, before making the spline grids finer with initialization from the coarser grid and continuing the training procedure. Grid extensions for all splines in a KAN can be performed independently and adaptively. We highlight that this technique is specific to the choice of splines as parameterizations of activation functions, and is particularly useful in the training process; see for example the discussion of spectral bias in the next section.
@@
We recollect the approximation theory established for compositionally smooth functions in [46].
@@ 3.2 Reparametrization of KANs and MLPs
Suppose that a function $f:[0,1]^{d}\rightarrow\mathbb{R}$ can be represented by a KAN with width $W\geq 1$, depth $L\geq 1$, and grid size $G$ with $k$-th order B-spline functions. Assume also that the weight $w_{b}=0$ in (3) for all of the activations $\phi_{l,i,j}$ appearing in the KAN. Then there exists an MLP with width $(G+2k+1)W$, depth at most $2L$, and activation function $\sigma_{k}=\max(0,x)^{k}$ which represents $f$.
We remark that the width in this result scales like $O(GW)$ so that the number of parameters in the MLP scales like $O(G^{2}W^{2}L)$, while the number of parameters in the KAN scales like $O(GW^{2}L)$. This appears to indicate that KANs with a large number of breakpoints at each node may be more efficient in expressing certain functions than MLPs, although we don’t know whether the preceding theorem is sharp.
Using these results, we can draw conclusions about the approximation capabilities of KANs by leveraging existing results about MLPs (see for instance [7, 38, 33, 64, 65, 27, 47, 61, 81, 63, 79, 78]). For example, we have the following result giving optimal approximation rates for very deep KANs on Sobolev spaces (see for instance [2] for the background on Sobolev spaces).
@@
This result, which follows immediately from Theorem 3.2 and the approximation rates for ReLU (and more generally piecewise polynomial) neural networks derived in [63], shows that very deep KANs attain an exceptionally good approximation rate on Sobolev spaces. In particular, in terms of the number of parameters $P$ they attain an approximation rate of $O(P^{-2s/d})$, while a classical (even non-linear) method of approximation can only attain a rate of $O(P^{-s/d})$ [21]. This phenomenon, which is often called superconvergence [20], also occurs for very deep ReLUk networks. However, it comes at the cost of parameters which are not encodable using a fixed number of bits and thus is not practically realizable [82, 63].
@@ 4 Spectral Bias
In this section, we study the spectral bias of KANs and compare it to that of MLPs. The spectral bias, or frequency principle, refers to the observation that neural networks trained with gradient descent tend to be biased toward lower frequencies, i.e. that lower frequencies are learned first. This phenomenon has been well-documented and studied for MLPs, see for instance [53, 84, 75, 76, 28, 51, 15, 9, 23, 86], and it is our purpose here to develop an analogous theory for KANs. We remark that although the spectral bias is considered a form of regularization which is desirable for machine learning tasks [53, 84, 51, 87, 23], for scientific computing applications it is typically important to capture all frequencies and so the spectral bias may negatively affect the performance of neural networks for such applications [71, 54, 15, 28].
@@ 4.1 Spectral bias theory for shallow KANs
We consider the spectral bias properties of KANs with a single layer. This theory is very similar to the theory developed in [28, 86, 56] for the spectral bias of single hidden layer MLPs. The key observation is that a single layer KAN is a linear model. In particular, we see that if $L=1$ then the KAN applied to an input $\mathbf{x}\in\mathbb{R}^{d}$ is (for simplicity, we consider the case of a KAN without the SiLu non-linearity)
@@
The convergence of gradient descent on the least squares regression is determined by the eigendecomposition of the Hessian matrix $M$ which is estimated in the following theorem. The theorem is essentially a generalization of the fact that the Gram matrix of the B-spline basis is well conditioned (see for instance [22], Theorem 4.2 in Chapter 5).
@@
Theorem 4.1 shows that away from $d^{\prime}(d-1)$ eigenvectors the matrix $M$ is well conditioned. This means that gradient descent will converge at the same rate in all directions orthogonal to these $d^{\prime}(d-1)$ eigenvectors. Note that since the number of eigenvectors we must remove is independent of the grid size $G$, we expect that when $G$ is relatively large most components of the KAN will converge at roughly the same rate toward the solution. Thus the KAN with a large number of grid points will not exhibit the same spectral bias toward low frequencies seen by MLPs. We remark that in constrast the Hessian associated with a two layer ReLU MLP with width $n$ has a condition number which scales like $n^{4}$ [28], which explains the strong spectral bias exhibited by ReLU MLPs.
@@
We note that the $d^{\prime}(d-1)$ eigenvectors which must be excluded is not an artifact of the proof. In fact, this is due to the fact that the KAN parameterization is not unique. Indeed, the constant function $f(x)=1$ can be parameterized in $d$ different ways by using the B-splines in each of the $d$ different directions. This ambiguity gives rise to directions in parameter space where the function parameterized by the KAN doesn’t change and this results in degenerate eigenvectors of the matrix $M$.
@@
In the first example, we take the same setting as in [53] and study the regression of a linear combination of waves of different frequencies. Consider the function prescribed as
@@
The phases $\varphi_{i}$ are uniformly sampled from $[0,2\pi]$ and we consider two cases of amplitudes: one with equal amplitude $A_{i}=1$ and another with increasing amplitude $A_{i}=0.1i$. We use a neural network, either ReLU MLP or KAN, to regress $f$ sampled at $200$ uniformly spaced points in $[0,1]$, with full batch ADAM iteration as the optimizer with a learning rate of $0.0003$. For MLPs, we train with $80000$ iterations as in [53]; for KANs, we only train with 8000 iterations. Normalized magnitudes of discrete Fourier transform at frequencies $k_{i}$ are computed as $|\tilde{f}_{k_{i}}/A_{i}|$ and averaged over 10 runs of different phases.
We plot the evolution of $|\tilde{f}_{k_{i}}/A_{i}|$ during training across all frequencies; see Figures 1, 2 for comparisons of MLPs and KANs with different sizes for equal and increasing amplitudes respectively. KANs suffer significantly less than MLPs from spectral biases. Once the size of KANs, especially the grid size and depth is large enough, KANs almost learn all frequencies at the same time, while even very deep and wide MLPs still have difficulties learning higher frequencies, even with $10$x epochs!

### v2

@@ Abstract
Kolmogorov-Arnold Networks (KAN) Liu et al. (2024e) were very recently proposed as a potential alternative to the prevalent architectural backbone of many deep learning models, the multi-layer perceptron (MLP). KANs have seen success in various tasks of AI for science, with their empirical efficiency and accuracy demonstrated in function regression, PDE solving, and many more scientific problems.
@@ 1 Introduction
Recently, in Liu et al. (2024e), a novel architecture called Kolmogorov-Arnold Networks (KANs) was proposed as a potentially more accurate and interpretable alternative to standard multi-layer perceptron (MLP) Cybenko (1989); Hornik et al. (1989). The KAN architecture leverages the Kolmogorov-Arnold representation theorem (KART) Kolmogorov (1956; 1957) to parameterize functions. KANs share the same fully connected structures as MLPs, while putting learnable activation functions on edges as opposed to a fixed activation function on nodes for MLPs. B-splines are used to parameterize the learned nonlinearity and in practice one can go beyond the two-layer construction indicated by KART. KANs can be conceptualized as a hybrid of splines and MLPs, and the combination of compositional nonlinearity with 1D splines contributes to both the accuracy and interpretability of KANs.
@@
In this article, we study the approximation theory and spectral bias of KANs and compare them with MLPs. Specifically, we show that any MLP with the ReLUk activation function can be reparameterized as a KAN with a comparable number of parameters. This establishes that the representation and approximation power of KANs is at least as great as that of MLPs. On the other hand, we also show that any KAN (without SiLU non-linearity) can be represented using an MLP. However, the number of parameters in the MLP representation is larger by a factor proportional to the grid size of the KAN. While we do not know if our construction is optimal in the deep case, this suggests that KANs with a large grid size may be more efficient at parameterizing certain classes of functions than MLPs. We combine these results with existing optimal approximation rates for MLPs to obtain approximation rates for KANs on Sobolev spaces.
@@
Next, we study the spectral bias phenomenon in the training of KANs. Standard MLPs with ReLU activations (or even tanh) are known to suffer from the spectral bias Rahaman et al. (2019); Xu et al. (2019a; b), in the sense that they will fit low-frequency components first. This is in contrast to traditional iterative numerical methods like the Jacobi method which learn high frequencies first Xu et al. (2019a). Although the spectral bias acts as a regularizer which improves performance for machine learning applications Rahaman et al. (2019); Zhang et al. (2021); Poggio et al. (2018); Zhang et al. (2020); Fridovich-Keil et al. (2022), for scientific computing applications it may be necessary to learn high-frequencies as well. To alleviate the spectral bias, high-frequency information has to be encoded using methods like Fourier feature mapping Sitzmann et al. (2020); Tancik et al. (2020); Benbarka et al. (2022); Novello et al. (2024), or one needs to use nonlinear activation functions more similar to traditional methods; see for example the hat activation function Hong et al. (2022) which resembles a finite element basis. We theoretically consider a single KAN layer and show that it does not suffer from the spectral bias by analyzing gradient descent for minimizing a least squares objective. Although our analysis is necessarily highly simplified, we argue that it provides some evidence and intuition showing that KANs will have a reduced spectral bias compared with MLPs.
@@
Finally, we study the spectral bias of KANs experimentally on a wide variety of problems, including 1D frequency fitting, fitting samples from a higher-dimensional Gaussian kernel, and solving the Poisson equation with a high-frequency solutions. Based upon the results of our experiments, KANs consistently suffer less from the spectral bias. This also helps explain why KANs are more subject to noise and overfitting, as observed in Shen et al. (2024), while grid coarsening (the opposite of grid extension) helps increase the spectral bias and reduces overfitting. Notice that the highest frequency that a single layer of KANs can learn is restricted to the number of grid points we use. The compositional structure associated with the depth also plays a role in learning the different frequencies. We study the effect of depths, widths, and grid sizes of KANs for the spectral bias, and draw conclusions about best practices when choosing KAN hyperparameters.
@@ 2 Prior Work
A theory on the approximation ability of KANs in the function class of compositionally smooth functions was proposed as KAT in Liu et al. (2024e). A convergence result independent of the dimension was obtained, leveraging the approximation theory of 1D splines. Compared with the universal approximation theorem of MLPs, KANs take advantage of the intrinsically low-dimensional compositional representation of underlying functions. This result shares an analogy to the rate in generalization error bounds of finite training samples, for a similar space studied for regression problems; see Horowitz & Mammen (2007); Kohler & Langer (2021), and also specifically for MLPs with ReLU activations Schmidt-Hieber (2020). On the other hand, for general Sobolev or Besov spaces, sharp approximation rates have been obtained for ReLU-MLPs (and more generally MLPs with most piecewise polynomial activation functions) Yarotsky (2017); Bartlett et al. (2019); Siegel (2023). These rates exhibit the curse of dimensionality, which is unavoidable due to the fact that Sobolev and Besov spaces with fixed smoothness are very large in high dimensions.
@@
There are also subsequent works exploring other parametrizations of activation functions in the KAN formulation, including special polynomials Aghaei (2024a); Seydi (2024a); SS (2024), rational functions Aghaei (2024b), radial basis function Li (2024); Ta (2024), Fourier series Xu et al. (2024a), and wavelets Bozorgasl & Chen (2024); Seydi (2024b). Active follow-up research focuses on applying KANs to various domains, such as partial differential equations Wang et al. (2024); Shukla et al. (2024); Rigas et al. (2024) and operator learning Abueidda et al. (2024); Shukla et al. (2024); Nehma & Tiwari (2024), graphs Bresson et al. (2024); De Carlo et al. (2024); Kiamari et al. (2024); Zhang & Zhang (2024), time series Vaca-Rubio et al. (2024); Genet & Inzirillo (2024b); Xu et al. (2024b); Genet & Inzirillo (2024a), computer vision Cheon (2024b); Azam & Akhtar (2024); Li et al. (2024a); Cheon (2024a); Seydi (2024b); Bodner et al. (2024), and various scientific problems Liu et al. (2024b; c); Yang et al. (2024); Herbozo Contreras et al. (2024); Kundu et al. (2024); Li et al. (2024b); Ahmed & Sifat (2024); Liu et al. (2024a); Peng et al. (2024); Pratyush et al. (2024). In the KAN 2.0 paper Liu et al. (2024d), multiplication was introduced as a built-in modularity of KANs and the connection between KANs and scientific problems was further established in terms of identifying important features, modular structures, and symbolic formulas. In this article, we focus on the original B-spline formulation, one of the reasons being its alignment with continual learning and adaptive learning; see also Rigas et al. (2024).
@@
The spectral bias of MLPs has been studied both experimentally and theoretically in a variety of works, see for instance Rahaman et al. (2019); Zhang et al. (2021); Xu et al. (2019a; b); Hong et al. (2022); Poggio et al. (2018); Cai & Xu (2019); Basri et al. (2020); Fridovich-Keil et al. (2022); Zhang et al. (2023); Ronen et al. (2019) and the references therein. This phenomenon is proposed as explaining the regularizing effect of stochastic gradient descent Rahaman et al. (2019); Fridovich-Keil et al. (2022). In Wang et al. (2024) an empirical analysis of the eigenvalues of the Neural Tangent Kernel Jacot et al. (2018) matrix is performed, which hints at the better performance of KANs for fitting high frequencies. However, the spectral bias of finite width KANs has not yet been considered, and it is this gap that we aim to close in this work.
@@ 3 Representation and Approximation
In this section, we study how to represent ReLUk MLPs using KANs with degree $k$ splines and vice-versa. Using this, we establish approximation rates for KANs from the corresponding results for MLPs. The measure of complexity that we consider in this section is the number of parameters of the KAN or MLP model. We remark that the number of parameters may not always be the most useful measure of complexity, especially considering that we are comparing different architectures, but it is generally indicative of the efficiency of the model. We leave the problem of comparing KANs and MLPs using different measures of complexity to future work.
@@ 3.1 Review of the KAN Architecture
We recall the following definitions of the Kolmogorov-Arnold Network (KAN) architecture introduced in Liu et al. (2024e). We refer to Liu et al. (2024e) for a thorough treatment of all details mentioned here.
@@
The Kolmogorov-Arnold representation theorem (KART) states that any multivariate continuous function on a bounded domain can be represented as a finite composition of univariate continuous functions and addition. Specifically for a continuous $f:[0,1]^{n}\to\mathbb{R}$, there exists continuous 1D functions $\phi_{q,p},\Phi_{q}$ such that (see for instance Braun & Griebel (2009))
@@
Inspired by KART corresponding to a two-layer neural network representation, the authors in Liu et al. (2024e) unified the inner and outer functions via the proposed KAN layers and generalized the idea of composition to arbitrary depths. A KAN can be represented by an integer array $[n_{0},n_{1},\cdots,n_{L}]$. Here $L$ is the number of layers and $n_{0}$ and $n_{L}$ are input and output dimensions respectively. $n_{l}$ denotes the number of neurons in the $l$-th neuron layer. Usually, we choose $n_{1}=\cdots=n_{L-1}=W$ as the width and refer to $L$ as the depth of the KAN. Layer $l$ maps a vector $\mathbf{x}_{l}\in\mathbb{R}^{n_{l}}$ to $\mathbf{x}_{l+1}\in\mathbb{R}^{n_{l+1}}$, via nonlinear activations and additions, in the matrix form as
@@
A very important structure of KANs proposed in Liu et al. (2024e) is the grid extension technique, inheriting the multi-level fine-graining from splines. One can first train KANs with fewer parameters to a desired accuracy, before making the spline grids finer with initialization from the coarser grid and continuing the training procedure. Grid extensions for all splines in a KAN can be performed independently and adaptively. We highlight that this technique is specific to the choice of splines as parameterizations of activation functions, and is particularly useful in the training process; see for example the discussion of spectral bias in the next section.
@@
We recollect the approximation theory established for compositionally smooth functions in Liu et al. (2024e).
@@ 3.2 Reparametrization of KANs and MLPs
Suppose that a function $f:[0,1]^{d}\rightarrow\mathbb{R}$ can be represented by a KAN with width $W\geq 1$, depth $L\geq 1$, and grid size $G$ with $k$-th order B-spline functions. Assume also that the weight $w_{b}=0$ in (3) for all of the activations $\phi_{l,i,j}$ appearing in the KAN. Then there exists an MLP with width $(G+2k+1)W^{2}$, depth at most $2L$, and activation function $\sigma_{k}=\max(0,x)^{k}$ which represents $f$.
We remark that the width in this result scales like $O(GW^{2})$ so that the number of parameters in the MLP scales like $O(G^{2}W^{4}L)$, while the number of parameters in the KAN scales like $O(GW^{2}L)$. This indicates that KANs with a large number of breakpoints at each node may be more efficient in expressing certain functions than MLPs. When the grids of the spline basis are different across each neuron pairs, which is natural after grid update in training, then there are $(G+2k+1)W^{2}$ different break points of the splines for each KAN layer. If we restrict ourselves to using only $1$ hidden layer of ReLUk for each KAN layer, as in the construction of the proof in appendix C, then by He et al. (2018) it follows that the width of ReLUk MLP has to be at least $(G+2k+1)W^{2}$ and the theorem is sharp. However, we don’t know whether the preceding theorem is sharp in the deep case.
Using these results, we can draw conclusions about the approximation capabilities of KANs by leveraging existing results about MLPs (see for instance Barron (1993); Leshno et al. (1993); Klusowski & Barron (2018); Siegel & Xu (2020; 2022); Hon & Yang (2022); Lu et al. (2021); Shen et al. (2022); Yarotsky (2018); Siegel (2023); Yang et al. (2023); Yang & Lu (2023)). For example, we have the following result giving optimal approximation rates for very deep KANs on Sobolev spaces (see for instance Adams & Fournier (2003) for the background on Sobolev spaces).
@@
This result, which follows immediately from Theorem 3.2 and the approximation rates for ReLU (and more generally piecewise polynomial) neural networks derived in Siegel (2023), shows that very deep KANs attain an exceptionally good approximation rate on Sobolev spaces. In particular, in terms of the number of parameters $P$ they attain an approximation rate of $O(P^{-2s/d})$, while a classical (even non-linear) method of approximation can only attain a rate of $O(P^{-s/d})$ DeVore (1998). This phenomenon, which is often called superconvergence DeVore et al. (2021), also occurs for very deep ReLUk networks. However, it comes at the cost of parameters which are not encodable using a fixed number of bits and thus is not practically realizable Yarotsky & Zhevnerchuk (2020); Siegel (2023).
@@ 4 Spectral Bias
In this section, we study the spectral bias of KANs and compare it to that of MLPs. The spectral bias, or frequency principle, refers to the observation that neural networks trained with gradient descent tend to be biased toward lower frequencies, i.e. that lower frequencies are learned first. This phenomenon has been well-documented and studied for MLPs, see for instance Rahaman et al. (2019); Zhang et al. (2021); Xu et al. (2019a; b); Hong et al. (2022); Poggio et al. (2018); Cai & Xu (2019); Basri et al. (2020); Fridovich-Keil et al. (2022); Zhang et al. (2023), and it is our purpose here to develop an analogous theory for KANs. We remark that although the spectral bias is considered a form of regularization which is desirable for machine learning tasks Rahaman et al. (2019); Zhang et al. (2021); Poggio et al. (2018); Zhang et al. (2020); Fridovich-Keil et al. (2022), for scientific computing applications it is typically important to capture all frequencies and so the spectral bias may negatively affect the performance of neural networks for such applications Wang et al. (2022); Rathore et al. (2024); Cai & Xu (2019); Hong et al. (2022).
@@ 4.1 Spectral bias theory for shallow KANs
We consider the spectral bias properties of KANs with a single layer. This theory is very similar to the theory developed in Hong et al. (2022); Zhang et al. (2023); Ronen et al. (2019) for the spectral bias of single hidden layer MLPs. The key observation is that a single layer KAN is a linear model. In particular, we see that if $L=1$ then the KAN applied to an input $\mathbf{x}\in\mathbb{R}^{d}$ is (for simplicity, we consider the case of a KAN without the SiLu non-linearity)
@@
The convergence of gradient descent on the least squares regression is determined by the eigen decomposition of the Hessian matrix $M$ which is estimated in the following theorem. The theorem is a generalization of the fact that the Gram matrix of the B-spline basis is well conditioned (see for instance DeVore & Lorentz (1993), Theorem 4.2 in Chapter 5).
@@
Spectral bias refers to the fact that lower frequencies converge much faster in training, which is related to the ill-condition property of the Hessian matrix in shallow models. Theorem 4.1 shows that away from $d^{\prime}(d-1)$ eigenvectors the matrix $M$ is well conditioned. This means that gradient descent will converge at the same rate in all directions orthogonal to these $d^{\prime}(d-1)$ eigenvectors. Note that since the number of eigenvectors we must remove is independent of the grid size $G$, we expect that when $G$ is relatively large most components of the KAN will converge at roughly the same rate toward the solution. Thus the KAN with a large number of grid points will not exhibit the same spectral bias toward low frequencies seen by MLPs.
@@
We remark that in constrast the Hessian associated with a two layer ReLU MLP with width $n$ has a condition number which scales like $n^{4}$ Hong et al. (2022), which explains the strong spectral bias exhibited by ReLU MLPs. On the other hand, as established in Remark 14 in Zhang et al. (2023), ReLUk MLP with width $n$ has a condition number at the order of $n^{2+2k}$. One can also observe in Sec 4.3 that they perform even worse for high-frequency tasks.
Remark 4.3.
We remark that in Wang et al. (2024) a spectral bias analysis for infinite width KANs is performed empirically using an analysis of the associated neural tangent kernel for specific tasks. Theorem 4.1 complements this work by giving an analysis in the finite width regime as well.
Remark 4.4.
We note that the $d^{\prime}(d-1)$ eigenvectors which must be excluded is not an artifact of the proof. In fact, this is due to the fact that the KAN parametrization is not unique. Indeed, the constant function $f(x)=1$ can be parameterized in $d$ different ways by using the B-splines in each of the $d$ different directions. This ambiguity gives rise to directions in parameter space where the function parameterized by the KAN doesn’t change and this results in degenerate eigenvectors of the matrix $M$.
@@
In the first example, we take the same setting as in Rahaman et al. (2019) and study the regression of a linear combination of waves of different frequencies. Consider the function prescribed as
@@
The phases $\varphi_{i}$ are uniformly sampled from $[0,2\pi]$ and we consider two cases of amplitudes: one with equal amplitude $A_{i}=1$ and another with increasing amplitude $A_{i}=0.1i$. We use a neural network, either ReLU MLP or KAN, to regress $f$ sampled at $200$ uniformly spaced points in $[0,1]$, with full batch ADAM iteration as the optimizer with a learning rate of $0.0003$. For MLPs, we train with $80000$ iterations as in Rahaman et al. (2019); for KANs, we only train with 8000 iterations. Normalized magnitudes of discrete Fourier transform at frequencies $k_{i}$ are computed as $|\tilde{f}_{k_{i}}/A_{i}|$ and averaged over 10 runs of different phases.
We plot the evolution of $|\tilde{f}_{k_{i}}/A_{i}|$ during training across all frequencies; see Figure 1 and Figure 7 in the appendix B for comparisons of MLPs and KANs with different sizes for equal and increasing amplitudes respectively. KANs suffer significantly less than MLPs from spectral biases. Once the size of KANs, especially the grid size and depth is large enough, KANs almost learn all frequencies at the same time, while even very deep and wide MLPs still have difficulties learning higher frequencies, even with $10$x epochs!